OpenStudy (anonymous):

how do i simplify:2x-2y/3x+3y muliplied by 6x+6y/x^2-y^2

OpenStudy (helder_edwin):

do you mean \[ \frac{2x-2y}{3x+3y}\cdot\frac{6x+6y}{x^2-y^2} \]

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

ok, so \[ \frac{2x-2y}{3x+3y}\cdot\frac{6x+6y}{x^2-y^2}= \frac{2(x-y)}{3(x+y)}\cdot\frac{6(x+y)}{(x-y)(x+y)} \] can you go from here?

OpenStudy (anonymous):

so would i just multiply the 2 x 6 first?

OpenStudy (helder_edwin):

no, first you simplify! which of factors go away?

OpenStudy (anonymous):

i think u would cancel (x+y) and (x-y) then u would simplify and get 4/(x+y)

OpenStudy (helder_edwin):

\[ \frac{2(x−y)}{3(x+y)}⋅\frac{6(x+y)}{(x−y)(x+y)}=\frac{2(x-y)}{3(x+y)}\cdot \frac{3\cdot 2(x+y)}{(x-y)(x+y)} \] do you get it?

OpenStudy (anonymous):

so when i simplify it i get this 6/(x-y) ?

OpenStudy (helder_edwin):

no the answer is \[ \frac{4}{x+y} \] do you get it?